You want the cover of a sports media guide to show two photos across it's width, the cover is 6.5 inches wide, and the left and right margins are 3.75 each. The space between the photos is .5 inches, how wide should you make the photos?
Write an equation to model the problem.

Answer by checkley77(12844)

6.5-(3.75-.5)=6.5-4.25=2.25 INCHES IS AVAILABLE FOR THE 2 PICTUIRES.
2.25/2=1.125 INCHES IS THE MAXIMUM WIDTH OF EACH PICTURE.